OpenStudy (anonymous):

is the following function a linear transformation? T:f(negative infinity,infinity) where T(F(x))=f(x+1)

OpenStudy (anonymous):

T(cf(x))=cf(x+1) =cT(f(x)) where c is a contant. is it corect?

OpenStudy (anonymous):

what are properties of linear transform ??

OpenStudy (anonymous):

i dont know myself

OpenStudy (anonymous):

T(cU)=cT(U) and T(u + v)=T(u)+T(V)

OpenStudy (anonymous):

we cant conclude T(cf(x))=cf(x+1)

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

our variable is x here not the hole thing...to my understanding the best we can do\[T(f(cx))=f(cx+1)=???\]

OpenStudy (anonymous):

sorry i have no idea @estudier

OpenStudy (anonymous):

thanks @mukushla

OpenStudy (anonymous):

@TuringTest ,any idea

OpenStudy (anonymous):

@satellite73 plz look at this

OpenStudy (turingtest):

@jacobian your analysis of the transformation obeying the scalar multiplication rule is correct the transformation is to add one to the argument

OpenStudy (turingtest):

\[T(f(x))=f(x+1)\]\[T(cf(x))=cf(x+1)\]\[cT(f(x))=cf(x+1)~~~~~\large\checkmark\]

OpenStudy (turingtest):

@jacobian still here?

OpenStudy (anonymous):

yah

OpenStudy (turingtest):

awesome for the next part you need to remember that the sum of any two polynomials is another polynomial, so we can always write\[f(x)+g(x)=h(x)\]so bear that in mind....

OpenStudy (turingtest):

so\[T[(f(x)+g(x)]=T(h(x))=h(x+1)\]\[=f(x+1)+g(x+1)=T(f(x))+T(g(x))~~~~\large\checkmark\]catch that?

OpenStudy (anonymous):

yes ,thanks a bulk

OpenStudy (turingtest):

anytime :)